# Benchmarking Zero-Shot Facial Emotion Annotation with Large Language Models: A Multi-Class and Multi-Frame Approach in DailyLife

### v1

@@ Abstract.
This study investigates the feasibility and performance of using large language models (LLMs) to automatically annotate human emotions in everyday scenarios. We conducted experiments on the DailyLife subset of the publicly available FERV39k dataset, employing the GPT-4o-mini model for rapid, zero-shot labeling of key frames extracted from video segments. Under a seven-class emotion taxonomy (“Angry,” “Disgust,” “Fear,” “Happy,” “Neutral,” “Sad,” “Surprise”), the LLM achieved an average precision of approximately 50%. In contrast, when limited to ternary emotion classification (negative/neutral/positive), the average precision increased to approximately 64%. Additionally, we explored a strategy that integrates multiple frames within 1-2 second video clips to enhance labeling performance and reduce costs. The results indicate that this approach can slightly improve annotation accuracy. Overall, our preliminary findings highlight the potential application of zero-shot LLMs in human facial emotion annotation tasks, offering new avenues for reducing labeling costs and broadening the applicability of LLMs in complex multimodal environments.
Annotation, large language model, gpt, zero-shot, image augmentation, scalable oversight, image sentiment analysis, benchmark
copyright: nonejournalyear: 2025ccs: Computing methodologies Computer graphicsccs: Computing methodologies Computer vision tasks
@@ 1. Introduction
Recent advancements in artificial intelligence (AI), particularly in the realm of large language models (LLMs), have opened new avenues for automating annotation tasks (Tan et al., 2024a). LLMs, such as Generative Pre-trained Transformer (GPT), possess sophisticated natural language understanding capabilities and operate effectively in zero-shot settings, where they can perform tasks without explicit prior training on specific datasets (Tan et al., 2024b). These models have demonstrated potential in various applications, from text generation to semantic understanding, suggesting their utility in assisting or even replacing human annotators (Zhang et al., 2024c; Thapa et al., 2023; Alizadeh et al., 2023).
@@
Furthermore, the latest iterations of LLMs integrate visual capabilities, enabling them to comprehend and interpret graphical information in conjunction with textual data (Zhao et al., 2024; Shao et al., 2024). This multimodal proficiency suggests that LLMs could become valuable tools for tasks that encompass both visual and linguistic components (Maniparambil et al., 2023; Hu et al., 2023; Bendeck and Stasko, 2025). By leveraging their ability to understand visual inputs and operate in zero-shot settings, LLMs have the potential to streamline the annotation process while maintaining both accuracy and efficiency (Jiang et al., 2024).
@@
Building on these capabilities, our study investigates the feasibility and performance of using LLMs for the automatic annotation of human emotions in everyday scenarios. Specifically, we employ the GPT-4o-mini model to conduct rapid, zero-shot labeling of key frames extracted from video segments within the DailyLife subset of the publicly available FERV39k dataset (Wang et al., 2022). Our experiments assess the model’s performance across two emotion taxonomies: a seven-class taxonomy encompassing “Angry,” “Disgust,” “Fear,” “Happy,” “Neutral,” “Sad,” and “Surprise,” and a ternary taxonomy categorizing emotions as negative, neutral, or positive.
@@
Our results indicate that the LLM attained an average precision of approximately 50% in the seven-class taxonomy, surpassing a simple baseline. This underscores the model’s ability to discern complex emotional states without task-specific training. Notably, when the classification was constrained to a simpler ternary classification, the average precision increased to around 64%, demonstrating the model’s enhanced performance in broader emotion categories. Additionally, we investigated strategies of integrating multiple frames within 1-2 second video clips to improve labeling performance and reduce annotation costs. This approach resulted a slight but notable improvement in accuracy.
@@
Our findings contribute to the growing body of research aimed at enhancing data annotation methods through AI. By examining the capabilities and limitations of zero-shot LLMs in emotion annotation tasks, we also discuss the potential for employing LLMs in practical applications to reduce costs and enhance scalability in real-world applications.
@@ 2.1. Emotion Annotation
2.2. LLM for Annotation
The emergence and application of LLMs have introduced unprecedented opportunities in the field of data annotation. An increasing number of researchers and practitioners have recognized the vast potential of LLMs for enhancing annotation processes (Mohta et al., 2023). As researchers continue to explore and leverage the advancing capabilities of LLMs, particularly in multimodal interactions (Zhang et al., 2024e) and improvements in processing power (Bender et al., 2021), the range of annotation tasks has expanded significantly. These tasks now encompass various data types, including text (Liu et al., 2024), audio (Du et al., 2024), images (Cheng et al., 2024; Sapkota et al., 2024), and specialized domain-specific data (Tang et al., 2024; Zhang et al., 2024d).
A recent survey has shed light on current trends and leading research in the application of LLMs for annotation tasks (Tan et al., 2024b). Within the scope of our study, which focuses on emotion annotation for image data, related work has explored various capabilities of LLMs. For instance, researchers have evaluated the ability of LLMs to predict emotions from captions generated from images-derived captions (Yang et al., 2023), perform image retrieval (Yang et al., 2024), and generate descriptive captions (Shvetsova et al., 2025). Notably, in early 2024, a study compared the performance of LLMs such as GPT-3.5, GPT-4, and Bard against traditional supervised models like Convolutional Neural Networks (CNNs) for emotion recognition in image data (Nadeem et al., 2024). The findings revealed that deep learning models specifically trained for this task generally achieved higher accuracy than LLMs.
However, despite the superior accuracy of traditional supervised models, they also present significant limitations. Nonetheless, LLMs offer the potential to achieve performance that is comparable to traditional models while reducing training and application costs. Therefore, in this study, we further optimized prompt engineering and reorganized annotation strategies to harness the capabilities and advantages of LLMs.
@@ 3.3.1. Zero-Shot Labeling
The annotation process was conducted using a zero-shot approach, wherein the GPT-4o-mini model was prompted with simple, predefined instructions to label the extracted key frames. No additional training or fine-tuning was performed on the model for the specific emotion annotation task. The prompts were designed to instruct the model to identify and classify the dominant emotion depicted in each frame based on visual and contextual cues.
To maximize cost efficiency, we utilized five specific frames from each video segment for annotation: the initial frame, the frame at the first quartile (Q1) position, the middle frame, the frame at the third quartile (Q3) position, and the last frame. This selection process reduces annotation counts while still capturing key emotional transitions within each segment. Subsequently, we applied different weighting strategies in the annotation process to derive comprehensive labels for the entire video segments based on these five selected frames. This approach balances the need for accurate emotion recognition with the practical constraints of annotation costs, ensuring that our methodology remains both effective and scalable.
@@ 3.3.2. Prompt Engineering
In our study, we implemented prompt engineering to effectively utilize the GPT-4o-mini model for emotion annotation in images. The prompt was meticulously crafted to guide the model’s responses by defining clear roles and providing specific instructions (Lin, 2024). Initially, a prompt was set to establish the model as a “professional image emotion analysis assistant,” explicitly listing the available emotion options derived from the predefined $EMOTION\_LABELS$. This foundational setup ensures that the model operates within the desired context and understands the classification framework. For each image (or multi-frame integrated image) to be analyzed, we constructed a user message that includes both textual instructions and the image itself, e.g., “This is an independent image frame, please analyze the emotion. Please analyze the emotion of the following image and select the most matching one from the above options, returning only the emotion name.
@@ 3.4.3. Annotation Strategy C1 (Seven-Class Taxonomy)
Strategy C1 is determining the predominant emotion by excluding the “Neutral” category. Specifically, if one emotion constitutes an absolute majority among the annotated frames after removing “Neutral,” that emotion is assigned to the video segment. However, if all five frames are labeled as “Neutral,” the segment is assigned the “Neutral” label. In cases where there is an equal distribution of different emotions, the emotion label of the middle frame is selected to represent the overall emotion of the video segment. This approach aims to enhance annotation accuracy by focusing on more distinctly positive or negative emotional states, thereby mitigating the ambiguous property of LLM in classifying the intermediate emotion of “neutral”.
@@ 3.4.4. Annotation Strategy D1 (Seven-Class Taxonomy)
Strategy D1 employs a multi-frame integration approach by concatenating the five selected frames into a single composite input. Specifically, the initial frame, Q1 position frame, middle frame, Q3 position frame, and final frame are sequentially joined to form a unified image input. This consolidated input is then presented to the GPT-4o-mini model for annotation in a single step.
@@ 3.4.5. Annotation Strategy A2 (Three-Class Taxonomy)
Strategy A2 adapts the results from Strategy A1 to the three-class emotion taxonomy. In this strategy, each of the five annotated frames from Strategy A1 is directly mapped to one of three broader categories: “Positive,” “Neutral,” or “Negative.” Specifically, emotions categorized as “Angry,” “Disgust,” “Fear,” and “Sad” are classified as “Negative,” while “Happy” and “Surprise” are classified as “Positive.” The “Neutral” labels are still “Neutral”.
@@ 3.4.7. Annotation Strategy C2 (Three-Class Taxonomy)
Strategy C2 involves first applying Strategy A2 to reorganize seven-class labels into three classes, followed by a strategy similar to Strategy C1 to mitigate the ambiguous property of LLM in classifying the intermediate emotion of “neutral”.
@@ 4.1. Evaluation Metrics
We assessed our annotation strategies using precision, recall, F1-score, support, and accuracy. Precision measures the proportion of correct predictions for each emotion, while recall evaluates the ability to identify all relevant instances. The F1-score balances precision and recall, making it useful for uneven class distributions. Accuracy reflects the overall correctness of the model. Additionally, we report macro average and weighted average to provide insights into performance across all classes, with macro average treating each class equally and weighted average accounting for class imbalance by weighting metrics based on class support. Support refers to the number of true instances for each emotion category in the dataset, providing context for the other metrics by indicating the distribution of classes.
@@ 4.2. Seven-Class Taxonomy
Overall, Strategies C1 and D1 demonstrate superior performance and cost-effectiveness in the demanding seven-class taxonomy tasks. Notably, Strategy D1 further reduces costs by minimizing the number of API requests and decreasing token lengths through preprocessing. This indicates that aggregation techniques and the integration of temporal context offer enhanced performance advantages in LLM’s zero-shot annotation tasks.
@@ 4.3. Three-Class Taxonomy
Overall, Strategies B2, C2, and D2 consistently outperformed Strategy A2, further highlighting the effectiveness of aggregation and integration methods in enhancing annotation accuracy within zero-shot classification approaches based on LLMs.
Furthermore, the macro average and weighted average metrics underscore the balanced performance of Strategies B2, C2, and D2 across all emotion categories. The macro average indicates that these strategies maintain consistent precision and recall across both common and rare classes, while the weighted average reflects their strong overall performance by accounting for the distribution of classes in the dataset.

### v2

# Zero-shot Emotion Annotation in Facial Images Using Large Multimodal Models: Benchmarking and Prospects for Multi-Class, Multi-Frame Approaches

@@ Abstract.
This study investigates the feasibility and performance of using large multimodal models (LMMs) to automatically annotate human emotions in everyday scenarios. We conducted experiments on the DailyLife subset of the publicly available FERV39k dataset, employing the GPT-4o-mini model for rapid, zero-shot labeling of key frames extracted from video segments. Under a seven-class emotion taxonomy (“Angry,” “Disgust,” “Fear,” “Happy,” “Neutral,” “Sad,” “Surprise”), the LMM achieved an average precision of approximately 50%. In contrast, when limited to ternary emotion classification (negative/neutral/positive), the average precision increased to approximately 64%. Additionally, we explored a strategy that integrates multiple frames within 1-2 second video clips to enhance labeling performance and reduce costs. The results indicate that this approach can slightly improve annotation accuracy. Overall, our preliminary findings highlight the potential application of zero-shot LMMs in human facial emotion annotation tasks, offering new avenues for reducing labeling costs and broadening the applicability of LMMs in complex multimodal environments.
Annotation, large language model, gpt, zero-shot, image augmentation, scalable oversight, image sentiment analysis
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 3rd International Workshop on Multimodal and Responsible Affective Computing; October 27–31, 2025; Dublin, Ireland.booktitle: Proceedings of the 3rd International Workshop on Multimodal and Responsible Affective Computing (MRAC ’25), October 27–31, 2025, Dublin, Irelandccs: Computing methodologies Computer graphicsccs: Computing methodologies Computer vision tasks
@@ 1. Introduction
Recent advancements in artificial intelligence (AI), particularly in the realm of large multimodal models (LMMs), have opened new avenues for automating annotation tasks (Zhang et al., 2025a; Tan et al., 2024a). LMMs, such as Generative Pre-trained Transformer (GPT), possess sophisticated natural language understanding capabilities and operate effectively in zero-shot settings, where they can perform tasks without explicit prior training on specific datasets (Tan et al., 2024b). These models have demonstrated potential in various applications, from text generation to semantic understanding, suggesting their utility in assisting or even replacing human annotators (Zhang et al., 2025b; Thapa et al., 2023; Alizadeh et al., 2023).
@@
Furthermore, the latest iterations of LMMs integrate visual capabilities, enabling them to comprehend and interpret graphical information in conjunction with textual data (Zhao et al., 2024; Shao et al., 2024). This multimodal proficiency suggests that LMMs could become valuable tools for tasks that encompass both visual and linguistic components (Maniparambil et al., 2023; Hu et al., 2023; Bendeck and Stasko, 2025). By leveraging their ability to understand visual inputs and operate in zero-shot settings, LMMs have the potential to streamline the annotation process while maintaining both accuracy and efficiency (Jiang et al., 2024). However, aside from capacity and performance considerations, the cost of computing resources is also very important. This matter is especially critical for individual users, students, small teams, and start-ups. Currently, because of the understandable business considerations, these groups rarely or even impossible to participate in the pricing process for computing resources such as API access and hardwares like graphic cards. Although we appreciate the discounts and grants offered to these groups, which show support from businesses and policies, these discounts do not reduce the overall use of computing resources. This means that their workflows and project goals remain limited by budget issues. In particular, some tasks or applications that depend heavily on computing resources may still not be easy to process.
@@
Building on these capabilities and considering ways to reduce input and output costs, our study investigates the feasibility and performance of using LMMs for the automatic annotation of human emotions in everyday scenarios. Specifically, we employ the GPT-4o-mini model to conduct rapid, zero-shot labeling of key frames extracted from video segments within the DailyLife subset of the publicly available FERV39k dataset (Wang et al., 2022). Our experiments assess the model’s performance across two emotion taxonomies: a seven-class taxonomy encompassing “Angry,” “Disgust,” “Fear,” “Happy,” “Neutral,” “Sad,” and “Surprise,” and a ternary taxonomy categorizing emotions as negative, neutral, or positive.
@@
Our results indicate that the LMM attained an average precision of approximately 50% in the seven-class taxonomy, surpassing a simple baseline. This underscores the model’s ability to discern complex emotional states without task-specific training. Notably, when the classification was constrained to a simpler ternary classification, the average precision increased to around 64%, demonstrating the model’s enhanced performance in broader emotion categories. Additionally, we investigated strategies of integrating multiple frames within 1-2 second video clips to improve labeling performance and reduce annotation costs. This approach resulted a slight but notable improvement in accuracy.
@@
Our findings contribute to the growing body of research aimed at enhancing data annotation methods through AI. By examining the capabilities and limitations of zero-shot LMMs in emotion annotation tasks, we also discuss the potential for employing LMMs in practical applications to reduce costs and enhance scalability in real-world applications.
@@ 2.1. Emotion Annotation
2.2. LMM for Annotation
The emergence and application of large language models (LLMs) and LMMs have introduced unprecedented opportunities in the field of data annotation. An increasing number of researchers and practitioners have recognized the vast potential of LLMs and LMMs for enhancing annotation processes (Mohta et al., 2023). As researchers continue to explore and leverage the advancing capabilities of LLMs, particularly in multimodal interactions (Zhang et al., 2024d) and improvements in processing power (Bender et al., 2021), the range of annotation tasks has expanded significantly. These tasks now encompass various data types, including text (Liu et al., 2024), audio (Du et al., 2024), images (Cheng et al., 2024; Sapkota et al., 2024), and specialized domain-specific data (Tang et al., 2024; Zhang et al., 2024c).
A recent survey has shed light on current trends and leading research in the application of LMMs for annotation tasks (Tan et al., 2024b). Within the scope of our study, which focuses on emotion annotation for image data, related work has explored various capabilities of LMMs. For instance, researchers have evaluated the ability of LMMs to predict emotions from captions generated from images-derived captions (Yang et al., 2023), perform image retrieval (Yang et al., 2024), and generate descriptive captions (Shvetsova et al., 2025). Notably, in early 2024, a study compared the performance of LMMs such as GPT-3.5, GPT-4, and Bard against traditional supervised models like Convolutional Neural Networks (CNNs) for emotion recognition in image data (Nadeem et al., 2024). The findings revealed that deep learning models specifically trained for this task generally achieved higher accuracy than LMMs.
However, despite the superior accuracy of traditional supervised models, they also present significant limitations. Nonetheless, LMMs offer the potential to achieve performance that is comparable to traditional models while reducing training and application costs. Therefore, in this study, we further optimized prompt engineering and reorganized annotation strategies to harness the capabilities and advantages of LMMs.
@@ 3.3.1. Zero-Shot Labeling
The annotation process used a zero-shot approach, prompting the GPT-4o-mini model with simple, predefined instructions to label extracted key frames. No additional training or fine-tuning was performed for the specific emotion annotation task. The prompts instructed the model to identify and classify the dominant emotion in each frame based on visual and contextual cues.
To optimize cost efficiency, five frames from each video segment were annotated: the initial frame, the first quartile (Q1), the middle frame, the third quartile (Q3), and the last frame. This sampling reduced annotation counts while capturing key emotional transitions. Different weighting strategies were then applied to derive comprehensive labels for each segment based on these frames. This approach balances accuracy in emotion recognition with cost constraints, ensuring the methodology remains effective and scalable. Apart from the predefined annotation strategy, we did not impose any additional human intervention on the LMM’s labeling process.
@@ 3.3.2. Prompt Engineering
In our study, we implemented prompt engineering to effectively utilize the GPT-4o-mini model for emotion annotation in images. The prompt was meticulously crafted to guide the model’s responses by defining clear roles and providing specific instructions (Lin, 2024). Initially, a prompt was set to establish the model as a “professional image emotion analysis assistant,” explicitly listing the available emotion options derived from the predefined EMOTION_LABELSEMOTION\_LABELS. This foundational setup ensures that the model operates within the desired context and understands the classification framework. For each image (or multi-frame integrated image) to be analyzed, we constructed a user message that includes both textual instructions and the image itself, e.g., “This is an independent image frame, please analyze the emotion. Please analyze the emotion of the following image and select the most matching one from the above options, returning only the emotion name.
@@ 3.4.3. Annotation Strategy C1 (Seven-Class Taxonomy)
Strategy C1 is determining the predominant emotion by excluding the “Neutral” category. Specifically, if one emotion constitutes an absolute majority among the annotated frames after removing “Neutral,” that emotion is assigned to the video segment. However, if all five frames are labeled as “Neutral,” the segment is assigned the “Neutral” label. In cases where there is an equal distribution of different emotions, the emotion label of the middle frame is selected to represent the overall emotion of the video segment. This approach aims to enhance annotation accuracy by focusing on more distinctly positive or negative emotional states, thereby mitigating the ambiguous property of LMM in classifying the intermediate emotion of “neutral”.
@@ 3.4.4. Annotation Strategy D1 (Seven-Class Taxonomy)
Strategy D1 employs a multi-frame integration approach by concatenating the five selected frames into a single composite input. Specifically, the initial frame, Q1 position frame, middle frame, Q3 position frame, and the final frames are sequentially joined along the temporal dimension, rather than across the channel, to form a unified image input. This consolidated input is then presented to the GPT-4o-mini model for annotation in a single step.
@@ 3.4.5. Annotation Strategy A2 (Three-Class Taxonomy)
Strategy A2 adapts the results from Strategy A1 to the three-class emotion taxonomy. In this strategy, each of the five annotated frames from Strategy A1 is directly mapped to one of three broader categories (Lee et al., 2023): “Positive,” “Neutral,” or “Negative.” Specifically, emotions categorized as “Angry,” “Disgust,” “Fear,” and “Sad” are classified as “Negative,” while “Happy” and “Surprise” are classified as “Positive.” The “Neutral” labels are still “Neutral”.
@@ 3.4.7. Annotation Strategy C2 (Three-Class Taxonomy)
Strategy C2 involves first applying Strategy A2 to reorganize seven-class labels into three classes, followed by a strategy similar to Strategy C1 to mitigate the ambiguous property of LMM in classifying the intermediate emotion of “neutral”.
@@ 4.1. Evaluation Metrics
We assessed our annotation strategies using precision, recall, F1-score, support, and accuracy (in Fig. 2). Precision measures the proportion of correct predictions for each emotion, while recall evaluates the ability to identify all relevant instances. The F1-score balances precision and recall, making it useful for uneven class distributions. Accuracy reflects the overall correctness of the model. Additionally, we report macro average and weighted average to provide insights into performance across all classes, with macro average treating each class equally and weighted average accounting for class imbalance by weighting metrics based on class support. Support refers to the number of true instances for each emotion category in the dataset, providing context for the other metrics by indicating the distribution of classes.
(a) Precision Comparison Across Strategies for Seven-Class Annotation: Individual Metrics and Averages. This graph illustrates the precision scores for each emotion category across all strategies, along with macro and weighted averages denoted by dashed and dash-dot lines, respectively.
(b) Precision Comparison Across Strategies for Three-Class Annotation: Individual Metrics and Averages. This graph highlights the precision scores for each emotion category (negative, neutral, positive) across all strategies, alongside macro and weighted averages.
Figure 2. Precision Comparison for Seven-Class and Three-Class Annotation Strategies. Both graphs showcase individual metrics and overall averages (macro and weighted) for each strategy.
@@ 4.2. Seven-Class Taxonomy
Source W df Q $p$-unc
Method 0.0129 2 60.37 $7.78\times 10^{-14}$
Table 2. Friedman Test Results for Merged Data (Strategy B1, C1, and D1). This table summarizes the outcome of a Friedman test applied to compare the prediction correctness across the merged dataset for three methods. The test indicates that there is a statistically significant difference in performance among the methods.
Correct_B1 Correct_C1 Correct_D1
Correct_B1 1.000000 1.000000 0.000218
Correct_C1 1.000000 1.000000 0.005641
Correct_D1 0.000218 0.005641 1.000000
Table 3. Dunn Post-Hoc Test Results (Bonferroni Adjusted) for Merged Data. Pairwise comparisons using Dunn’s test reveal that while two of the Strategies (B1 and C1) do not differ significantly, the Strategy D1 shows a statistically significant difference from Strategy B1 and C1.
To further validate these observations, a non-parametric Friedman test was conducted on the merged dataset (comprising the results from Strategies B1, C1, and D1). As shown in Table 2, the Friedman test revealed a statistically significant difference among these strategies (p¡0.05). This overall significance suggests that at least one strategy performs differently compared to the others.
Subsequently, pairwise comparisons were performed using Dunn’s post-hoc test with Bonferroni correction (see Table 3). The results indicate that there is no statistically significant difference between Strategies B1 and C1. In contrast, Strategy D1 is statistically significantly different from both B1 and C1 (p¡0.05). These statistical findings provide robust evidence that the multi-frame integration approach (Strategy D1), which harnesses temporal context, exhibits a distinct performance profile compared to the majority voting methods.
Furthermore, when considering the macro average and weighted average metrics, both Strategies C1 and D1 demonstrated improved balanced performance across all emotion categories. The macro average reflects consistency in performance across less frequent categories, while the weighted average accounts for class imbalances, further reinforcing the overall effectiveness of these strategies.
@@ 4.3. Three-Class Taxonomy
Table 5. Friedman Test Results for Merged Data (Methods B2, C2, and D2).
Source W df Q $p$-unc
Method 0.014 2 67.178 $2.585\times 10^{-15}$
Table 6. Dunn Post-Hoc Test Results (Bonferroni Adjusted) for Merged Data.
Correct_B2 Correct_C2 Correct_D2
Correct_B2 1.000000 0.000195 0.000251
Correct_C2 0.000195 1.000000 1.000000
Correct_D2 0.000251 1.000000 1.000000
Overall, Strategies B2, C2, and D2 consistently outperformed Strategy A2, further highlighting the effectiveness of aggregation and integration methods in enhancing annotation accuracy within zero-shot classification approaches based on LMMs. Furthermore, the macro average and weighted average metrics underscore the balanced performance of Strategies B2, C2, and D2 across all emotion categories. The macro average indicates that these strategies maintain consistent precision and recall across both common and rare classes, while the weighted average reflects their strong overall performance by accounting for the distribution of classes in the dataset.
To further validate these observations, a non-parametric Friedman test was conducted on the merged dataset (comprising the results of Strategies B2, C2, and D2). As summarized in Table 5, the Friedman test revealed a statistically significant difference among the three methods (p ¡ 0.05), which supports the notion that at least one strategy produces a performance profile that is distinct from the others. Subsequent pairwise comparisons using Dunn’s post-hoc test with Bonferroni adjustment (see Table 6) demonstrated that Strategy B2 significantly differs from both Strategies C2 and D2 (p ¡ 0.05), while no significant difference was found between Strategies C2 and D2. This statistical evidence reinforces our conclusion that the choice of aggregation strategy plays a critical role in determining annotation accuracy and consistency.
In summary, the combined performance metrics and the statistical tests confirm that the aggregation and integration methods (especially Strategies C2 and D2) not only yield superior overall accuracy but also provide a more balanced performance across diverse emotion categories, thereby underscoring the importance of temporal context and strategic data integration in zero-shot annotation tasks.